OpenStudy (anonymous):

Find the value of y when x equals 1 in this equation. 2x + (-7y) = -54

OpenStudy (anonymous):

So if we make x = 1 what will the equation look like?

OpenStudy (anonymous):

idek

OpenStudy (anonymous):

2x + (-7y)=-54 2x - 7y =-54 2x1 -7y=-54 2 -7y =-54 -7y=-56 y=8